# You Name It, I Run It: An LLM Agent to Execute Tests of Arbitrary Projects

### v1

@@ 2.3.5. Web Search
For popular projects, there is often information available on the web that can help with the installation process and that goes beyond the information available in the repository itself. To retrieve such information, ExecutionAgent queries a search engine with the following query “How to install the ¡LANGUAGE¿ project ’¡PROJECT¿’ from source code?”, where ¡LANGUAGE¿ and ¡PROJECT¿ are replaced with the main programming language and the name of the project, respectively. The approach then extracts all text from the top-five results and asks an LLM to summarize it into the same structure as shown in Figure 3. In case a search result does not contain any relevant installation hints, the model is instructed to return “Irrelevant web page”.
2.4. Feedback Loop
Based on the prompt ingredients obtained in the preparation phase, the second phase of our approach is a feedback loop that iteratively invokes tools to install the project and execute its test suite. As shown in lines 8 to 31 of Algorithm 1, the feedback loop consists of two nested loops. The inner loop runs a series of commands, guided by the LLM, until either the test suite completes successfully or a configurable command limit (default: 40) is reached. The outer loop allows for multiple attempts to install and execute the test suite, with a configurable maximum number of attempts (default: 3). This design allows recovery from occasional failures where the installation process encounters an error, preventing the tests from running. By allowing multiple attempts, the system can adjust its approach to overcome such errors. At the start of each outer loop iteration, the system incorporates lessons learned from the previous attempt (if any) into the command prompt. These lessons are generated by prompting the LLM to analyze the previous sequence of thoughts and commands, identify challenges encountered, and suggest adjustments for the next iteration (line 30). Each iteration of the inner loop consists of three steps, as outlined in Figure 1 and detailed in the following.
2.4.1. Step 1: LLM Agent to Select the Next Command
The core of our approach is an LLM that selects the next command to execute based on the current state of the installation process. To guide the LLM in selecting the next command, the approach constructs a command prompt that contains the prompt ingredients obtained in the preparation phase, as well as a summary of the commands executed so far. The command prompt consists of static sections, which are the same any time ExecutionAgent is used, and dynamic sections, which are either created based on the preparation phase (line 12) or updated at the end of each iteration of the inner loop (line 26). Specifically, the prompt contains the following sections:
@@ 2.4.1. Step 1: LLM Agent to Select the Next Command
    Agent role (static): Defines the primary task of the agent and the success criteria. Concretely, the agent’s role is to install the given project and ensure that the test cases execute properly.
@@
    • Write scripts to install the project and execute its tests.
@@
    • Analyze the test execution results and summarize them by providing the number of executed, passed, and failed tests.
@@
    Installation hints (dynamic): Contains the installation hints obtained from any existing CI/CD scripts and the web search. It also contains the previous attempts lessons (outer loop feedback).
@@
    Instruction for creating the next command prompt (static): Specifies that the LLM should provide the next command to execute based on the context available in the prompt. We define the expected format of the LLM’s response in a TypeScript interface, as shown in Figure 4, so the LLM can answer in JSON compatible with the given interface. The expected format includes the agent’s thoughts, as well as the name of the tool to invoke and the arguments to pass to the tool. The rationale for asking the LLM to describe its thoughts is two-fold: First, it has been shown empirically to improve the quality of the LLM’s responses (Wei et al., 2022). Second, it allows for easier debugging and understanding of the LLM’s decisions.
Expected format for selecting the next command interface Response {// Express your thoughts based on the information that you have collected so far, the possible steps that you could do next, and also your reasoning. thoughts: string; tool: {name: string; args: Record<string, any>; }; } Here is an example of a command call that you can output: { "thoughts": "I need to check the files and folders available within this repository.", "tool": {"name": "linux_terminal", "args": {"cmd": "ls"} } }
@@ 2.4.2. Step 2: Invoking Tools
To enable our approach to take the steps necessary for installing the project and running its tests, we provide four tools to the agent. ExecutionAgent invokes these tools based on the LLM’s response to the command prompt (lines 17 to 22).
@@
Similar to human developers, access to a terminal is crucial to successfully set up a project and run its tests, e.g., to install dependencies, configure the environment, or list available files. We provide the agent with the capability to execute any command available in a Linux terminal via the terminal tool. The tool takes a command to execute as input and returns the output of the command.
@@
While the terminal would, in principle, be sufficient to perform any kind of operation on the system, we provide two additional tools for interacting with the file system. These tools are meant to emulate the developer’s interaction with text editors, where they open a file, read parts of it, and sometimes make changes by writing to it. The read_file tool takes a file path as input and returns the content of the file. The write_file tool takes a file path and content as input and writes the content to the file. The latter tools is particularly useful for writing the scripts expected as the output of ExecutionAgent, e.g., a Dockerfile that creates a container and a install.sh script that installs the project and runs its tests.
@@
In addition to the tools described above, we provide the agent with a special tool called task_done. This tool is used to signal that the agent has successfully completed its task, i.e., that the project has been installed and its tests have been executed. The tool takes a natural language description that explains why the agent has finished the task as its argument.
@@ 2.4.3. Step 3: Summarization and Extraction
The output of tools can be verbose and contain lots of irrelevant information. For example, suppose the agent executes a command to list all files in a directory or reads the content of a file, then simply returning the output of the tool to the LLM for the next command would quickly fill up the available prompt size. Even with LLMs that offer a large prompt size, it is beneficial to reduce the amount of information in the prompt to keep the agent focused on the most important information and to reduce the overall costs of the approach.
@@ 2.4.4. Control Center
The control center combines the three steps described above and manages the interaction between the LLMs and the tools. In particular, it performs the following tasks:
  • Parse the LLM output and validate whether it conforms to the specified output format. We empirically observe that the LLM output is syntactically well-formed in almost all cases.
  • Invoke the next step as specified in Algorithm 1, e.g., by invoking a specific command.
  • For any tool invocation, check whether the command is returning within a configurable timeout (default: five minutes). In case the command has not terminated within this timeout, the control center provides the so far produced output to the LLM agent and asks to decide between three options: wait another five minutes, provide some input to the running tool, or kill the command. Providing some input to the already running tool is useful for interactive tools that require user input, e.g., a user confirming with “y” that some package should be installed.
  • Clean the output of the tools by removing terminal color characters and other special characters, e.g., printed to display a progress bar.
When the agent selects the task_done command, it indicates that the agent believes that the project has been successfully installed and that its test suite has been executed. Before terminating ExecutionAgent, the control center validates that the agent has indeed met all goals. This involves checking whether three files exist: a file to create a container (e.g., a Dockerfile), an installation script to be executed within the container, and a file that contains the test results (i.e., the number of passed, failed, and skipped tests). If all goals are met, ExecutionAgent returns the installation and test scripts to the user (line 21). Otherwise, the control center provides feedback to the LLM agent, pointing out what exactly is missing, and asks for a new command.
Figure 5 gives an example of the scripts output by ExecutionAgent for the OpenVPN project. The two scripts show the configuration chosen by ExecutionAgent, the installed dependencies, and the steps to install and execute the tests of the project.
@@ 3. Evaluation
  • RQ3 (ablation study): What is the impact of different components and configurations of ExecutionAgent?
@@ 3.1.1. Implementation and Model
ExecutionAgent is implemented in Python and bash. To allow for an isolated execution of ExecutionAgent itself, we use a Docker container. The agent is empowered by OpenAI’s GPT-4o-mini model, which we access through their Python API.
@@ 3.1.2. Metrics
The task performed by ExecutionAgent consists of two sub-tasks: (i) building and/or installing the project, and (ii) running the test suite. We measure how effective the approach is at performing these tasks by measuring the successful build rate, i.e., the proportion of projects where the approach manages to correctly build and install the project, and the successful testing rate, i.e., the proportion of projects where the approach manages to run the test suite. To determine these rates, we manually inspect the produced scripts and the test results.
@@ 3.1.3. Dataset
We gather a set of 50 open-source projects from GitHub that provide three properties: (i) The projects should cover different programming languages. To this end, we sample for each of the following languages ten projects that have that language as their dominant language or second dominant language: Python, Java, C, C++, and JavaScript. Because many projects contain code in multiple languages, the dataset overall covers 14 programming languages.
(ii) There exists a ground truth of test execution results that we can extract from the logs on a CI/CD platform. As different projects use different platforms, we collect ground truth data from GitHub Actions, CircleCI, Jenkins, CirrusCI, and occasionally, a project’s website (e.g, tensorflow). (iii) Each project has at least 100 starts and 100 commits with at least one commit in the past 6 months (prior to collection). Table 1 provides an overview of the projects in the dataset. Projects in our dataset have an median number of 10K commits, 45K stars, and 1.4K test cases.
@@
pytest P 3762 3640 11 111 3762 3757 5 0
react JS, TS, Rust 5509 5509 0 0 9736 0 0 0
@@ 3.1.4. Baselines
Given that LLMs have seen large amounts of data, including documentation on how to install projects, we leverage this feature to ask an LLM to generate a script that installs and runs test cases of an arbitrary project in a specific programming language. The prompt we use is given in Figure 6. As our dataset is gathered by focusing on five main languages, we generate one specialized script for each of these five languages. Following that, each project is attempted to be installed using the script corresponding to its main language.
@@
Instead of creating a specialized agent for the task of automatically setting up arbitrary projects, one could also use a general-purpose LLM-based agent. We compare against such an agent, AutoGPT [^2] https://github.com/Significant-Gravitas/AutoGPT (Yang et al., 2023), which given a task description, autonomously reasons about a task, makes a plan, executes, and updates the plan over multiple iterations. Similar to ExecutionAgent, AutoGPT may call tools, such as web search, reading and writing files, and executing Python code. As a task description, we provide the input shown in Figure 7. We use the same model and provide the same budget (max number of iterations) to AutoGPT as for ExecutionAgent.
@@ 3.2. Effectiveness
The results of applying ExecutionAgent to the 50 projects are reported in Table 1 and summarized in Table 2. ExecutionAgent successfully installs and tests 33 out of 50 projects. Out of the 33 successfully tested projects, ExecutionAgent achieves results identical to the ground truth in 17/33, while the rest (16 out 33) have an average deviation of 15.4%. In Table 2, we consider ExecutionAgent results to be close to the ground truth if the average deviation is less than 10%. Overall, these results show that ExecutionAgent is effective at executing the test suites of a large number of projects, and that the results are close to or equal to the ground truth in most cases.
@@
Comparison with LLM scripts
In Table 2, we compare ExecutionAgent to the three baselines. The general-purpose, LLM-generated scripts successfully build many projects (29/50), but then often often fail to execute the test suites (only 5/29 succeed). Inspecting the results, we find that the LLM scripts often do not account for the differences between the setup required by a regular user and the development setup. The development setup often involves additional steps, such as installing additional software (e.g., a compiler, glibc-32, or a testing framework), recompiling some dependencies, or changing configuration files. In contrast, the usage setup is much simpler, often using an already complete requirement/setup file, such as setup.py for Python and pom.xml for Java. Because the LLM scripts do not account for these differences, they often fail to execute the test suites, whereas ExecutionAgent is able to iteratively fix unexpected errors.
Comparison with AutoGPT
Even though AutoGPT is given the same budget as ExecutionAgent, it builds only 9/50 projects and successfully runs the tests suites for only four of them. The results show that the task of running the tests of arbitrary projects is non-trivial, and that a specialized agent, such as ExecutionAgent is more effective at such tasks. Fundamentally, ExecutionAgent differs from AutoGPT by using meta-prompting, by managing the memory of already performed commands more effectively, and by using more sophisticated and robust tools.
Comparison with Flapy
The right-most column of Table 2 compares with the Flapy baseline. Because this baseline only targets Python, we apply it only to the ten projects in our dataset that have Python as their main language. The results show that Flapy reaches results similar to LLM-generated, general-purpose scripts in terms of building projects. However, like the LLM scripts, it also struggles to execute the test suites, and cannot successfully complete the tests of any of the ten projects. Given the same ten Python projects, ExecutionAgent successfully builds and tests six of them, all with testing results identical to the ground truth.
@@
Results close to ground truth 29 / 50 4 / 50 2 / 50 0 / 10

### v2

@@ 2.3.5. Web Search
Some projects maintain explicit and structured documentation on external platforms, such as readthedocs.com, which can provide clearer installation instructions than their respective GitHub repositories. Additionally, search engines efficiently retrieve the most relevant resources based on a few keywords, reducing the need for exhaustive exploration.
To leverage web resources, ExecutionAgent queries a search engine with “How to install the ¡LANGUAGE¿ project ’¡PROJECT¿’ from source code?”, where ¡LANGUAGE¿ and ¡PROJECT¿ correspond to the project’s primary programming language and name, respectively. Instead of parsing webpages sequentially—which would increase the number of queries and associated costs—the approach extracts text from the top five results and asks an LLM to summarize them. The summarization process extracts relevant build and test dependencies, as well as installation steps, including commands and configuration details. If a search result lacks useful installation hints, the model is instructed to return: “This page does not contain setup-related information.”
2.4. Feedback Phase
Based on the prompt ingredients obtained in the preparation phase, the second phase of our approach is a feedback phase that iteratively invokes tools to build the project and execute its test suite. As shown in lines 8 to 31 of Algorithm 1, the feedback phase consists of two nested loops. The inner loop runs a series of commands, guided by the LLM, until either the test suite completes successfully or a configurable command limit (default: 40) is reached. The outer loop allows for multiple attempts to build and execute the test suite, with a configurable maximum number of attempts (default: 3). This design allows recovery from occasional failures where the build process encounters an error, preventing the tests from running. By allowing multiple attempts, the system can adjust its approach to overcome such errors. At the start of each outer loop iteration, the system incorporates lessons learned from the previous attempt (if any) into the command prompt. These lessons are generated by prompting the LLM to analyze the previous sequence of thoughts and commands, identify challenges encountered, and suggest adjustments for the next iteration (line 30). Each iteration of the inner loop consists of three steps, as outlined in Figure 1 and detailed in the following.
2.4.1. Step 1: LLM Selects the Next Command
The core of our approach is an LLM that selects the next command to execute based on the current state of the installation process. To guide the LLM in selecting the next command, the approach constructs a command prompt that contains the prompt ingredients obtained in the preparation phase, as well as a summary of the commands executed so far. The command prompt consists of predefined static sections, and dynamic sections, which are either created based on the preparation phase (line 12) or updated at the end of each iteration of the inner loop (line 26). Specifically, the prompt contains the following sections:
@@ 2.4.1. Step 1: LLM Selects the Next Command
    Agent role (static): Defines the primary task of the agent and the success criteria. Concretely, the agent’s role is to build the given project and ensure that the test cases execute properly.
@@
    • Write scripts to build the project and execute its tests.
@@
    • Analyze the test execution results and summarize them by providing the number of passing, failing, and skipped tests.
@@
    Installation hints (dynamic): Contains the installation hints obtained from any existing CI/CD scripts and the web search. It also contains lessons learned from previous iterations of the outer feedback loop.
@@
    Instruction for creating the next command prompt (static): Specifies that the LLM should provide the next command to execute based on the context available in the prompt. We define the expected format of the LLM’s response in a TypeScript interface, as shown in Figure 4. The expected format includes the agent’s thoughts, as well as the name of the tool to invoke and the arguments to pass to the tool. The rationale for asking the LLM to describe its thoughts is two-fold: First, it has been shown empirically to improve the quality of the LLM’s responses (Wei et al., 2022). Second, it allows for easier debugging and understanding of the LLM’s decisions. If the LLM fails to respond in this format, the control center returns an error message to the LLM, mentioning the usage of the wrong format and reminding the LLM of the correct one.
Test results interface Response {// Express your thoughts based on the information that you have collected so far, the possible steps that you could do next, and also your reasoning. thoughts: string; tool: {name: string; args: Record<string, any>; }; } Here is an example of a command call that you can output: { "thoughts": "I need to check the files and folders available within this repository.", "tool": {"name": "linux_terminal", "args": {"cmd": "ls"} } }
@@ 2.4.2. Step 2: Invoking Tools
To enable our approach to take the steps necessary for building the project and running its tests, we provide four tools to the agent. ExecutionAgent invokes these tools based on the LLM’s response to the command prompt (lines 17 to 22).
@@
Similar to human developers, access to a terminal is crucial to successfully set up a project and run its tests, e.g., to build dependencies or list available files. We provide the agent with the capability to execute any command available in a Linux terminal via the linux_terminal tool. The tool takes a command to execute as input and returns the output of the command.
@@
While the terminal would, in principle, be sufficient to perform any kind of operation on the system, we provide two additional tools for interacting with the file system. These tools are meant to emulate the developer’s interaction with text editors, where they open a file, read parts of it, and sometimes make changes by writing to it. The read_file tool takes a file path as input and returns the content of the file. The write_file tool takes a file path and content as input and writes the content to the file. The latter tool is particularly useful for writing the scripts expected as the output of ExecutionAgent, e.g., a Dockerfile that creates a container and a install.sh script with the sequence of commands to build dependencies, compile or build the project, and eventually run the test suite.
@@
In addition to the tools described above, we provide the agent with a special tool called task_done. This tool is used to signal that the agent has successfully completed its task, i.e., that the project has been built and its tests have been executed. The tool expects a natural language description that explains why the agent has finished the task.
@@ 2.4.3. Step 3: Summarization and Extraction
The output of tools can be verbose and contain lots of irrelevant information. For example, suppose the agent executes a command to list all files in a directory or reads the content of a file, then simply returning the output of the tool to the LLM for the next command would quickly exceed the available prompt size. Even with LLMs that offer a large prompt size, it is beneficial to reduce the amount of information in the prompt to keep the agent focused on the most important information and to reduce the overall costs of the approach.
@@ 2.4.4. Control Center
The control center orchestrates the three steps described above, managing the interaction between the LLMs and the tools. Specifically, it performs the following tasks:
  • Parse the LLM output and validate whether it conforms to the specified output format.
  • Invoke the next step as specified in Algorithm 1, e.g., by executing a specific command.
  • Monitor tool execution by tracking launched processes within the container and detecting when they terminate. Since some commands may take a long time before producing any output, merely watching stdout and stderr is insufficient. Instead, the control center ensures that process completion is explicitly detected. The complete stdout and stderr outputs are returned to the LLM once the process terminates.
  • Enforce a configurable timeout (default: five minutes) for any command execution. If a command does not finish within this interval, the control center provides the LLM agent with the available output and process status. The LLM then decides among three options: (1) wait for another five minutes, (2) provide input to the running tool (useful for interactive prompts, such as confirming an installation with “y”), or (3) terminate the command.
  • Clean the output of tools by removing terminal color codes and other special characters, such as those used for progress bars.
When the agent selects the task_done command, it indicates that the agent believes that the project has been successfully built and that its test suite has been executed. Before terminating ExecutionAgent, the control center validates that the agent has indeed met all goals. This involves checking whether three files exist: a file to create a container (e.g., a Dockerfile), an installation script to be executed within the container, and a file that contains the test results (i.e., the number of passed, failed, and skipped tests). If all goals are met, ExecutionAgent returns the installation and test scripts to the user (line 21). Otherwise, the control center provides feedback to the LLM agent, pointing out what exactly is missing, and asks for a new command. Figure 5 gives an example of the scripts produced for the OpenVPN project. The two scripts show the configuration chosen by ExecutionAgent, the installed dependencies, and the steps to build the project and execute its tests.
@@ 3. Evaluation
  • RQ3 (ablation study): What is the impact of different components and configurations?
@@ 3.1.1. Implementation and Model
ExecutionAgent is implemented in Python and bash. To allow for an isolated execution of ExecutionAgent itself, we use a Docker container. The agent is based on OpenAI’s GPT-4o-mini model, which we access through their Python API. The usage of other models is possible as long as they support JSON output format specified through an interface. At the moment of writing the paper, Claude, DeepSeek, and Llama models support this feature.
@@ 3.1.2. Metrics
The task performed by ExecutionAgent consists of two sub-tasks: (i) building the project, and (ii) running the test suite. We measure how effective the approach is at performing these tasks by measuring the successful build rate, i.e., the proportion of projects where the approach manages to correctly build the project, and the successful testing rate, i.e., the proportion of projects where the approach manages to run the test suite. To determine these rates, we manually inspect the produced scripts and the test results.
@@ 3.1.3. Dataset
To construct a dataset for evaluating ExecutionAgent, we gather 50 open-source projects from GitHub. The following describes the selection criteria and process.
Criteria for selecting projects
We select projects that meet the following criteria: (i) Diversity in programming languages: We select ten projects each for Python, Java, C, C++, and JavaScript, ensuring that each project has one of these languages as either the dominant or second most-used language. (ii) Availability of ground truth test execution results: We ensure that test execution results are accessible through logs from CI/CD platforms. As different projects use different platforms, we collect ground truth data from GitHub Actions, CircleCI, Jenkins, CirrusCI, and occasionally, from a project’s website (e.g., for TensorFlow). Our selection process prioritizes projects that maintain publicly accessible, detailed reports of test execution results, such as build/test logs on CI/CD platforms. One indicator of this is the presence of a CI/CD badge in the repository’s README. (iii) Project maturity and activity: Each selected project has at least 100 stars and 100 commits, with at least one commit made within the six months prior to data collection. This criterion ensures that the repositories are active and maintained.
Dataset construction process
Our dataset construction follows a systematic process to ensure quality and relevance. We search for the most popular projects for each of the five primary languages and rank them based on their popularity (number of stars and forks). We then manually inspect projects, starting at the top of the ranked list, to check whether they meet the selection criteria described above. If a project does not meet the criteria, we move to the next project in the list. We continue this process until we have collected 50 projects that meet the criteria.
Dataset statistics
The projects in our dataset exhibit significant activity and complexity. The median project has approximately 10K commits, 45K stars, and 1.4K test cases. The code in the 50 selected projects is written in 14 programming languages. When considering only test code, the projects cover 8 languages, including Kotlin, Scala, and TypeScript, in addition to the five primary languages. On average, 87% of the tests are written in the main language of the project. Some projects write test cases in different languages than their main language, though. For example, the CPython project has many test cases written in Python that test code implemented in C.
From another perspective, 38 out of the 50 projects use automated build tools or provide configuration files for them.
All projects in the dataset are capable of running on a Linux system, as confirmed by logs retrieved from CI/CD results. The dataset also highlights variations in CI/CD usage: 33 out of 50 projects contain scripts for both building and testing, 10 projects have build scripts only, and 7 projects lack both build and test scripts. Additionally, 9 projects contain submodules, indicating a higher degree of structural complexity. Table 1 lists all 50 projects and their characteristics.
@@
pytest P 3762 3640 11 111 3805 3775 11 119
react JS, TS, Rust 5509 5509 0 0 9764 9738 1 25
@@ 3.1.4. Baselines
LLMs have seen large amounts of data, including documentation on how to install projects. We leverage this feature to ask an LLM to generate a script that prepares an environment to build an arbitrary project and run its tests. As our dataset is gathered by focusing on five main languages, we ask the LLM to create one specialized script for each of these five languages. The prompt we use is given in Figure 6. Given the script created in response to this prompt, we attempt to build a project using the script corresponding to the project’s main language.
@@
Instead of creating a specialized agent for the task of automatically setting up arbitrary projects, one could also use a general-purpose LLM-based agent. We compare against such an agent, AutoGPT [^2] https://github.com/Significant-Gravitas/AutoGPT (Yang et al., 2023), which, given a task description, autonomously reasons about a task, makes a plan, executes, and updates the plan over multiple iterations. Similar to ExecutionAgent, AutoGPT may call tools, such as web search, reading and writing files, and executing Python code. As a task description, we provide the input shown in Figure 7. We use the same model and provide the same budget (maximum number of iterations) to AutoGPT as for ExecutionAgent.
@@ 3.2. Effectiveness
The results of applying ExecutionAgent to the 50 projects are reported in Table 1 and summarized in Table 2. ExecutionAgent successfully builds and tests 33 out of 50 projects. Out of the 33 successfully tested projects, ExecutionAgent achieves results identical to the ground truth in 17/33, while the rest (16 out 33) have an average deviation of 15.4%. In Table 2, we consider ExecutionAgent results to be close to the ground truth if the average deviation is less than 10%. Overall, these results show that ExecutionAgent is effective at executing the test suites of a large number of projects, and that the results are close to or equal to the ground truth in most cases.
@@
Another perspective that we analyze is the presence of configurations for automated build tools in the project repositories. Among the 50 projects, 38 use automated build tools, such as Make and Cmake (in 14 projects), Npm and Yarn (13 projects), and Maven/Gradle (9 projects). ExecutionAgent successfully runs the tests of 28 out of the 38 projects that use automated build tools, while the approach succeeds for 5 out of the 12 projects not using such tools. On the one hand, this result shows that the configuration and automation available within the project itself is a factor in the success of ExecutionAgent. On the other side, the result also shows that ExecutionAgent is capable of handling projects that lack such automation.
We also study the presence and impact of submodules. Submodules usually are standalone repositories that are included within the target project as a dependency and should be installed along with the target project. Our dataset has 9 projects that use submodules. We find that ExecutionAgent struggles with such projects, as only 2 out of 9 reach a successful test suite execution.
@@
Ground truth deviation ¡ 10% 29 / 50 4 / 50 2 / 50 0 / 10
Finally, we compare the individual test case outcomes between ExecutionAgent and the ground truth. In 17 out of 33 tested projects, the same test sets were executed with matching outcomes. Among the remaining 16 projects, 99.9% of passing test cases aligned with the ground truth, but deviation arose primarily in failing test cases. A detailed analysis of five projects with deviations revealed three key factors contributing to these differences. First, some projects are structured into independent modules with separate builds and test suites. For example, in the Activiti project, a failure in compiling the bpmn converter module halted the remaining testing process, though an existing option could have allowed the continuation of tests in other modules, which ExecutionAgent did not invoke. Second, incomplete dependency installations may cause test failures or errors. In the Axios project, for instance, the absence of a headless Chromium binary resulted in an error, preventing 8 test cases from executing. Third, we attribute certain differences to test flakiness as errors occurred without clear indications of missing dependencies or setup issues.
Comparison with LLM scripts
In Table 2, we compare ExecutionAgent to the three baselines. The general-purpose, LLM-generated scripts successfully build many projects (29/50), but then often often fail to execute the test suites (only 5/29 succeed). Inspecting the results, we find that the LLM scripts often do not account for the differences between the setup required by a regular user and the development setup. The development setup often involves additional steps, such as installing additional software (e.g., a compiler, glibc-32, or a testing framework), recompiling some dependencies, or changing configuration files. In contrast, the usage setup is much simpler, often using an already complete requirement/setup file, such as setup.py for Python and pom.xml for Java. Because the LLM scripts do not account for these differences, they often fail to execute the test suite, whereas ExecutionAgent is able to iteratively fix unexpected errors.
Comparison with AutoGPT
Even though AutoGPT is given the same budget as ExecutionAgent, it builds only 9/50 projects and successfully runs the tests suites for only 4 of them. The results show that the task of running the tests of arbitrary projects is non-trivial, and that a specialized agent, such as ExecutionAgent is more effective at such tasks. Fundamentally, ExecutionAgent differs from AutoGPT by using meta-prompting, by managing the memory of already performed commands more effectively, and by using more sophisticated and robust tools.
Comparison with Flapy
The right-most column of Table 2 compares with the Flapy baseline. Because this baseline only targets Python, we apply it only to the 10 projects in our dataset that have Python as their main language. The results show that Flapy reaches results similar to LLM-generated, general-purpose scripts in terms of building projects. However, like the LLM scripts, it also struggles to execute the test suites, and cannot successfully complete the tests of any of the 10 projects. Given the same 10 Python projects, ExecutionAgent successfully builds and tests 6 of them, all with testing results identical to the ground truth.